OpenStudy (flvs.net):

Armando read 20 pages of his 200-page book in 4 hours. At this rate, how long will it take him to read the entire book? 20 hours 40 hours 4 hours 5 hours

OpenStudy (alexalas):

so 20/4=?

OpenStudy (flvs.net):

5

OpenStudy (alexalas):

so that is 5 pages per hour correct?

OpenStudy (flvs.net):

Awesome!! can you help me with more please?

OpenStudy (alexalas):

sure

OpenStudy (flvs.net):

Marisa painted one-half of her bedroom in three-fourths of an hour. At this rate, how long would it take her the paint the entire room? three-eighths of an hour eight-thirds of an hour four-sixths of an hour six-fourths of an hour

OpenStudy (alexalas):

3/4 +3/4=?

OpenStudy (flvs.net):

1.5

OpenStudy (alexalas):

yes but keep it in fraction form

OpenStudy (flvs.net):

how do i convert it??

OpenStudy (anonymous):

Quick question what did you get for the first question you asked?

OpenStudy (flvs.net):

5

OpenStudy (alexalas):

yes but keep it in fraction form

OpenStudy (anonymous):

5 hours?

OpenStudy (flvs.net):

yea

OpenStudy (anonymous):

thats wrong

OpenStudy (alexalas):

5 pages for 1 hour

OpenStudy (flvs.net):

how do i convert it into a fraction?

OpenStudy (flvs.net):

how is it wrong? @Emma_1997

OpenStudy (alexalas):

1.5 to fraction put a 1 under it so 1.5/1 then multiply by 4 so (1.5*4)/(1*4)=6/4

OpenStudy (alexalas):

not five hours for one page but 5 pages for one hour

OpenStudy (flvs.net):

so it's 6/4? but wait it's improper sooo...

OpenStudy (alexalas):

the answer is 6/4. in word form that is (six-fourths) and that is the answer

OpenStudy (anonymous):

the first question is wrong because \[20 * 10 = 200 \] so you would do \[4 * 10 =?\] to get your answer for how many hours it would take her to read the entire book

OpenStudy (flvs.net):

okay thank you, i have two more can you help?

OpenStudy (flvs.net):

so 40 pages

OpenStudy (anonymous):

yess

OpenStudy (alexalas):

no it takes 40 hours

OpenStudy (anonymous):

no 40 hours

OpenStudy (flvs.net):

that's what i meant and THANK YOU GUYS, can you help me with more>

OpenStudy (alexalas):

oh I am sorry i mislead you 5/1 is the ratio for pages/hours so 5/1=x/200 you cross multiply and you get 40 hours per 200 pages

OpenStudy (flvs.net):

it's fine but ii have two more.

OpenStudy (alexalas):

lets see the first one

OpenStudy (flvs.net):

Margie calculated that she would spend $175 on school supplies this year. She actually spent $97.50 on school supplies. What is Margie’s percent of error? 77.3% 79.5% 44.2% 4.4%

OpenStudy (alexalas):

umm i am a little rusty with this one give me a few minutes

OpenStudy (flvs.net):

okay, hbu @Emma_1997

OpenStudy (anonymous):

give me a second im a little rusty too

OpenStudy (flvs.net):

okay, i suck at 7th grade math

OpenStudy (alexalas):

alright 97.50-175=?

OpenStudy (flvs.net):

-77.5

OpenStudy (alexalas):

so 77.5/175=?

OpenStudy (flvs.net):

.44285714__________________________________ it's repeated

OpenStudy (alexalas):

move the decimal over 2 places to make percentage so .442=44.2%

OpenStudy (flvs.net):

thank you!! >_< last one Maxine spent 15 hours doing her homework last week. This week she spent 18 hours doing homework. She says that she spent 120% more time doing homework this week. Is she correct? Show your work to justify your decision.

OpenStudy (alexalas):

change 120% to decimal form

OpenStudy (alexalas):

or divide 18 by 15

OpenStudy (flvs.net):

1.2

OpenStudy (alexalas):

so 15*1.2=?

OpenStudy (flvs.net):

18

OpenStudy (alexalas):

so is 120% of 15 18?

OpenStudy (flvs.net):

yes?

OpenStudy (alexalas):

alright there are two ways of doing this one 18/15=1.2=120% or 120%=1.2, 1.2*15=18 using either way yields the correct answer it depends on what is easier.

OpenStudy (alexalas):

if you need some more help go to https://www.mathsisfun.com/ it explains a lot of things

OpenStudy (flvs.net):

i missed one out of em'

OpenStudy (alexalas):

which one?

OpenStudy (flvs.net):

Margie calculated that she would spend $175 on school supplies this year. She actually spent $97.50 on school supplies. What is Margie’s percent of error? 77.3% 79.5% 44.2% 4.4% this one

OpenStudy (alexalas):

was it not 44.2%

OpenStudy (flvs.net):

no

OpenStudy (alexalas):

what was the answer?

OpenStudy (flvs.net):

idk

OpenStudy (alexalas):

no 44.2% is the answer

OpenStudy (flvs.net):

can you give me a medal?

OpenStudy (flvs.net):

thx

OpenStudy (flvs.net):

you get one too